# A Universal Scheme for Dynamic Partitioned Shortest Path Index

### v2

@@ 6.2. Experiment Result Analysis
Exp 2: Effect of Partition Method. Next, we further evaluate the effect of the partition method on the PSP indexes (PCH, PTD, and PPLL). In particular, we test PCH and PPLL with Bubble, PUNCH, KaHyPar, and METIS on road networks (NY and FL) while only testing PTD with PUNCH and Bubble by referring to Exp 1. Figure 11 shows that PUNCH and KaHyPar perform better than other partition methods in most cases due to smaller overall border number $|B|$ and average partition border number $\overline{B_{i}}$, which indicates that $|B|$ and $\overline{B_{i}}$ are important factors in the partition method selection. We select PUNCH as the default planar partition method due to its superior PSP index performance. It is worth noting that we also test PPLL with KaHyPar, METIS, and SCOTCH on complex networks. However, it fails to complete the index construction on the smallest graph GO within 48 hours due to the large and condensed overlay graph, which reveals that the PSP index with planar partition is generally unsuitable for complex networks.
Exp 3: Effect of Partition Number. As shown in Figure 12 (a)-(b), the indexing time of PCH and PTD increased as $k$ increased from 4 to 256. This is because a larger $k$ could lead to a larger overlay graph that cannot leverage thread parallelization, thus resulting in higher indexing time. By contrast, PPLL’s indexing time first decreases and then increases in NY while tending to decrease in FL because a larger $k$ could also lead to a smaller $\overline{|B_{i}|}$, thus reducing the overlay graph construction time. In Figure 12 (c)-(d), we observe that a smaller $k$ generally improves the query time but leads to worse index update efficiency. Therefore, it is crucial to balance the query and update efficiency by selecting the appropriate partition number. We set the default value of $k$ as 32 since it provided satisfactory performance in most cases.
Exp 4: Comparison with Non-partitioned SP Algorithms. We conduct a comparison between PSP indexes (including PCH, PTD, PPLL, and CT) and their non-partitioned counterparts (namely CH (Geisberger et al., 2008), TD (Ouyang et al., 2018), PLL (Akiba et al., 2013)). As depicted in Figure 13, PSP indexes generally provide faster indexing time, smaller index size, and better scalability than their non-partitioned counterparts. For instance, PCH and PTD exhibit up to 2.76×2.76$\times$2.76 × and 2.25×2.25$\times$2.25 × faster index construction compared to CH and TD, respectively. Moreover, CT demonstrates 72.3×72.3$\times$72.3 × and 3.58×3.58$\times$3.58 × improvement over PLL in terms of indexing time and index size, respectively. However, PSP indexes may compromise query processing efficiency and index update time. It is worth noting that it also shows that planar structure is more suited for road networks, while core-periphery structure is better suited for complex networks, as PPLL does not outperform PLL in terms of indexing time and scalability.
Exp 5: Effectiveness of PSP Strategy. As shown in Figure 14 (a)-(d), the no-boundary strategy has the minimum index construction (about $14-35\%$ speed-up) and index update time (up to two orders of magnitude speed-up compared with the pre-boundary strategy for PTD and PCH), which indicates the superiority of our no-boundary strategy. Suffering from the highest index construction time, the post-boundary strategy has the same query time as the pre-boundary strategy but a significant update efficiency improvement (14.636.7×14.6-36.7$\times$14.6 - 36.7 × speed up for PTD). To further demonstrate the effectiveness of our post-boundary strategy, we test the same-partition query efficiency as presented in Figure 14 (e)-(f). The post-boundary archives about 3.76.3×3.7-6.3$\times$3.7 - 6.3 × and 4.830.5×4.8-30.5$\times$4.8 - 30.5 × speed-up compared with the no-boundary strategy for PTD and PPLL, respectively. Such a significant performance gain comes from the post-boundary’s correct local partition index. Finally, we evaluate the effectiveness of pruning-based overlay graph optimization by measuring the average vertex degree of the overlay graph, as shown in Table 3. We test PUNCH and HEP on road networks while the KaHyPar and HEP on complex networks. The degree decreases with the pruning-based optimization in most cases. It cannot improve PUNCH since PUNCH already has a great partition result.
@@
Figure 15. Comparison of All Shortest Path Indexes. Bar: Left, Ball: Right
@@
Exp 6: Comparison of All PSP Algorithms. We compare our proposed PSP indexes with the existing PSP indexes in Figure 15. We first analyze the performance of our proposed indexes against their counterparts: 1) P-TD-P has the same query performance and index size as FHL, but it is much faster to update with $1$ 202.7 × speed-up. 2) N-CH-P has the fastest index construction and maintenance efficiency and the smallest index size among all planar PSP indexes, achieving up to 3.4×3.4$\times$3.4 ×, 1004×1004$\times$1004 ×, and 7.6×7.6$\times$7.6 × speed-up over FHL. 3) N-PC-CP has faster index construction and update than CT since it utilizes faster CH for tree index. P-PT-CP outperforms CT in terms of query processing as it is faster in same-partition queries. 4) N-TS-HP has faster index construction, smaller index size, faster query and update than G-Tree by leveraging TD for overlay index construction. 5) Sketch can scale up to very large graphs (the only algorithm that can be applied on graph FR), but its query efficiency is much slower than N-PC-CP and P-PT-CP due to the light index.
@@ 6.3. PSP Application Guidance
Based on the experimental analysis in Section 6.2, we clarify the misconceptions mentioned in Section 1 and conclude the following observations: 1) all PSP indexes are essentially interrelated because they all consist of three dimensions in our PSP index scheme: graph partition method, PSP strategy, and shortest path algorithm. Moreover, we could combine the elements of these three dimensions to create a new PSP index; 2) it is important to choose the right partition structure for different types of graphs. Generally, for road networks, planar and hierarchical PSP indexes are more suitable than core-periphery PSP due to their efficient query and index update efficiency. On the other hand, core-periphery PSP indexes are more suitable for complex networks due to their excellent scalability; 3) although PSP indexes typically have faster indexing efficiency, smaller index size, and better scalability than their non-partitioned counterparts, they sacrifice query and index update efficiency.
@@
Figure 16. Our Decision Tree of PSP Index Design
7. Related Work
Recently, multiple surveys about the SP computation have emerged, focusing on either specific graph types (like transportation networks (Bast et al., 2016) or complex networks (Zhang et al., 2022)) or static graphs (Madkour et al., 2017; Li et al., 2017; Wu et al., 2012). (Zhang et al., 2021a) investigates the computation in dynamic graphs, but it does not consider the partitioned counterpart. The partitioned pathfinding framework is proposed in (Chondrogiannis and Gamper, 2014, 2016), however, the dynamic version is not discussed and they can be covered by the traditional PSP strategy. In this paper, we study the dynamic PSP problem systematically and explore its performance in both road networks and complex networks by involving the state-of-the-art discovery of SP algorithms and partition methods. Next, we discuss and categorize the existing PSP methods based on our scheme:
1) Pre-Boundary + Search/All-Pair + Hierarchy/Planar: These methods (HiTi (Jung and Pramanik, 2002), Graph Separators, Customizable Route Planning (Delling et al., 2017) (Holzer et al., 2009), ParDiSP (Chondrogiannis and Gamper, 2014, 2016)) pre-compute the shortest distance between boundaries (some hierarchically) to guide the search. In a broad sense, Arc-flag (Möhring et al., 2007) and SHARC (Bauer and Delling, 2010) also belong to this category. G-tree (Zhong et al., 2013; Li et al., 2019a) uses dynamic programming to compute the distance between layer’s all-pair information to replace searching, and it is widely used in kNN (Zhong et al., 2015), ride-sharing (Ta et al., 2017), time-dependent routing (Wang et al., 2019), and machine learning-based path finding (Huang et al., 2021). They are slow to construct due to Pre-B and slower to query to direct search;
2) Pre-Boundary+Search/PLL+Planar: COLA (Wang et al., 2016) builds the labels for the skyline shortest path on the overlay graph to answer the constrained shortest path query. This structure’s construction suffers from Pre-B and query suffers from searching;
3) Pre-Boundary+TD/TD+Planar: FHL (Liu et al., 2021, 2022b, 2022a) partition builds the TD both within and between partitions for multi-dimensional skyline paths. As validated in Exp 4, our Post-B can speedup Pre-B’s construction dramatically;
4) Pre-Boundary+PLL/ PLL+Planar: T2Hop (Li et al., 2019c, 2020b) utilizes two layers of PLL to reduce the complexity of long range time-dependent paths, and this structure’s performance is limited by PLL;
5) No-Boundary+ PLL/TD+Core: Core-Tree (Li et al., 2020a; Zheng et al., 2022) is the baseline that we already discussed;
6) Pre-Boundary+Search/All-Pair/PLL+Sketch: This category works on huge graphs where index is nearly impossible so only a small number of landmarks are selected to either help prune the search (Das Sarma et al., 2010; Wang et al., 2021; Farhan et al., 2022) or approximate result (Gubichev et al., 2010; Qiao et al., 2012).
In sum, all the existing PSP methods have their limitation, warranting the comprehensiveinvestigation embarked on in this paper.
@@ 8. Conclusions
In this work, we systematically study the partitioned shortest path index and propose a universal PSP scheme by concomitantly considering three dimensions: PSP strategy, SP algorithm, and partition structure. For the PSP strategy, we propose two new strategies and pruned-boundary optimization for better index performance. We also provide index maintenance solutions for both traditional and novel PSP strategies. For partition structure, we propose a new path-oriented partition method classification. Furthermore, we construct our PSP index scheme by seamlessly coupling those three dimensions. We also propose a PSP index generator and five new PSP indexes that are either more efficient in query or update than the current state-of-the-art. Systematic and extensive experimental evaluations fully validate our findings and prove the effectiveness of our new indexes guided by the insightful PSP index scheme. Finally, PSP index design guidelines are provided to help future researchers and practitioners design their indexes. Our future work will address recommendation aspects for such designs.

### v3

@@ 7.2. Experiment Result Analysis
Exp 2: Effect of Partition Method on PSP Index. Next, we choose Bubble, PUNCH, KaHyPar, and METIS as the partition methods according to Exp 1 and evaluate their effect on PSP indexes. We test PCH and PTD on road networks NY, FL, W, and US since they are infeasible on complex networks with large treewidth. We test PPLL on road networks and complex network GO and do not present the result on US since PPLL cannot finish index construction within 24 hours for any partition methods. Figure 11 shows that PUNCH outperforms other partition methods in most cases due to smaller border numbers $|B|$ and $\overline{B_{i}}$ as evidenced in Exp 1, which reveals that they are the most important factors in the partition method selection. For complex networks that PUNCH cannot be applied, KaHyPar has the best performance. We have tried to test PPLL on larger complex networks. However, it fails to complete the test within 24 hours due to the large and condensed overlay graph, which reveals that the planar PSP index is generally unsuitable for complex networks. Therefore, we select PUNCH and KaHyPar as the default planar partition methods for road networks and complex networks, respectively. Besides, we also evaluate the effect of partition number as presented in the appendix (Zhang et al., [n. d.]) due to limited space. In short, we set the default partition number as 32 for all datasets since it provided satisfactory performance in most cases.
Table 4. Effectiveness of Overlay Optimization
PSP Index PCH PTD PPLL
Dataset Method $\tilde{D}$ $t_{c}$ $t_{q}$ $t_{u}$ $t_{c}$ $t_{q}$ $t_{u}$ $t_{c}$ $t_{q}$ $t_{u}$
NY Bubble w/o opt 187.45 112.8 18.04 0.20 134.5 18.06 0.42 70.27 1743 11.77
w/ opt 163.76 104.2 16.90 0.19 128.0 17.85 0.38 68.03 1561 10.00
PUNCH w/o opt 53.53 4.67 3.29 0.02 5.33 0.32 0.016 25.94 9.70 0.19
w/ opt 52.68 4.65 3.21 0.01 5.28 0.30 0.02 25.87 9.56 0.18
FL Bubble w/o opt 235.80 143.6 44.56 0.06 195.9 33.75 0.12 748.7 3032 1.28
w/ opt 211.87 138.0 38.82 0.05 187.8 31.37 0.09 742.4 2921 0.98
PUNCH w/o opt 37.42 3.57 13.98 0.03 5.69 0.17 0.05 368.7 4.83 0.24
w/ opt 37.26 3.43 13.81 0.03 5.61 0.16 0.04 365.7 4.34 0.24
$\tilde{D}$: average vertex degree of $\tilde{G}$; $t_{c}$: indexing time (s); $t_{q}$: query time (ms); $t_{u}$: index update time (s).
Figure 12. Effectiveness of PSP Strategy. Bar: Left, Ball: Right
Exp 3: Effectiveness of Overlay Graph Optimization. We evaluate the effectiveness of overlay graph optimization by reporting $t_{c}$, $t_{q}$, $t_{u}$, and average degree $\tilde{D}$ of $\tilde{G}$ on NY and FL. We do not use large networks since they fail to report the results of PCH, PTD, and PPLL simultaneously. As shown in Table 4, it enhances the index performance in most cases with evident improvement in Bubble-based partitions (achieving $14.6\%$, $8.3\%$, $14.7\%$, and $17.8\%$ speed-up on $\tilde{D}$, $t_{c}$, $t_{q}$, and $t_{u}$). The improvement on PUNCH is marginal since it already has a great partition result. Nevertheless, this optimization improves the index performance by reducing the overlay graph density with a theoretical guarantee of the index correctness and we will use it by default in the following experiments.
Figure 13. Performance of Different Index Update Types. Bar: Left, Ball: Right
Figure 14. Edge Weight Batch Updates of PTD. Bar: Left, Polyline: Right
Figure 15. Performance of Different Query Types
Exp 4: Effectiveness of PSP Strategy. We evaluate the effectiveness of the proposed PSP strategies on PCH, PTD, PPLL, and CT in Figure 12. The no-boundary always the fastest in update, achieving up to 185.2×185.2$\times$185.2 ×, 91.3×91.3$\times$91.3 ×, 1759.4×1759.4$\times$1759.4 ×, and 5358.8×5358.8$\times$5358.8 × faster than the pre-boundary for PCH, PTD, PPLL, and CT. Moreover, it also outperforms the pre-boundary in index construction time in PCH and PTD with up to 2.4×2.4$\times$2.4 × and 1.7×1.7$\times$1.7 × faster. This is because no-boundary replaces the time-consuming Dijkstra’s searches among the boundary vertices with index query. Nevertheless, it is slow in query since the same-partition queries require concatenating the overlay and partition indexes. The post-boundary addresses this drawback by repairing the partition indexes, achieving equivalent query time as pre-boundary. Moreover, since it needs to repair all-boundary-pair distances, it is time-consuming when the query efficiency of $\tilde{L}$ is low. For example, because the CH query processing is relatively slow, the index construction and maintenance of PCH with post-boundary is even slower than the pre-boundary as it involves twice all-pair searches among the boundaries. Nevertheless, the post-boundary could achieve a significant update efficiency improvement (14.659.3×14.6-59.3$\times$14.6 - 59.3 × for PTD) for PTD, PPLL, and CT that use 2-hop labeling (TD and PLL) as the underlying SP index.
Exp 5: Evaluation of Update Types. As shown in Figure 13, no-boundary is much faster (up to 185.2×185.2$\times$185.2 ×, 91.3×91.3$\times$91.3 ×, 1759.4×1759.4$\times$1759.4 × for PCH, PTD, and PPLL) than pre-boundary in all update types. The post-boundary works better than the pre-boundary on PTD and PPLL since it replaces all-pair searches with index query. Note that the performance of edge/vertex insertion (deletion) is almost equivalent to edge decrease (increase) since this structural update essentially shares a similar update mechanism with the edge weight update. Furthermore, the batch update for PTD is tested by varying batch size from 100 to 500. As shown in Figure 14, PTD-No and PTD-Post always outperform PTD-Pre. As the batch size increases, the improvement decreases since the all-pair boundary searches in the pre-boundary could be shared by different edge updates. Nonetheless, PTD-No and PTD-Post also share the computation for different edge updates during the all-pair boundary distance check and outperform PTD-Pre.
Exp 6: Performance of Different Query Types. We classify the queries into four types based on $s$ and $t$ locations: Q1 (same partition), Q2 (both boundary), Q3 (one boundary one non-boundary), Q4 (different partitions). As shown in Figure 15, for PTD and PPLL, Q2 is the fastest since it only uses $\tilde{L}$. By contrast, Q3 and Q4 are much slower due to distance concatenation. For Q1, the post-boundary is the same with pre-boundary since they both only use the partition index, while the no-boundary has to concatenate $\tilde{L}$ and $\{L_{i}\}$. Lastly, PCH has similar query efficiency on different query types and strategies due to its search manner.
@@
Figure 16. Comparison with Non-partitioned SP Indexes. (Bar: Left axis, Ball: Right axis)
Figure 17. PSP Indexes Comparison. Bar: Left, Ball: Right
Exp 7: Comparison with Non-partitioned SP Algorithms. We conduct a comparison between PSP indexes (including PCH, PTD, PPLL, and CT) and their non-partitioned counterparts (namely CH (Geisberger et al., 2008), TD (Ouyang et al., 2018), PLL (Akiba et al., 2013)). As depicted in Figure 16, PSP indexes generally provide faster indexing time, smaller index size, and better scalability than their non-partitioned counterparts. For instance, PCH and PTD exhibit up to 2.76×2.76$\times$2.76 × and 2.25×2.25$\times$2.25 × faster index construction compared to CH and TD, respectively. Moreover, CT demonstrates 72.3×72.3$\times$72.3 × and 3.58×3.58$\times$3.58 × improvement over PLL in terms of indexing time and index size, respectively. However, PSP indexes generally compromise query processing efficiency and even index update time (except for PTD). Besides, it can also be seen that planar structure is more suited for road networks, while core-periphery structure is better suited for complex networks, as only CT can finish tests on complex networks while PPLL cannot.
@@
Exp 8: Comparison of All PSP Algorithms. We compare our proposed PSP indexes with the existing ones in Figure 17. We first analyze the performance of our proposed indexes: 1) P-TD-P has the same query efficiency and index size as FHL but is much faster to update with $1$ 202.7 × speed-up; 2) N-CH-P has the fastest index construction and maintenance efficiency and the smallest index size among all planar PSP indexes, achieving up to 3.4×3.4$\times$3.4 ×, 1004×1004$\times$1004 ×, and 7.6×7.6$\times$7.6 × speed-up over FHL; 3) N-PC-CP has faster index construction and update than CT since it utilizes faster CH for tree index. P-PT-CP outperforms CT in terms of query processing as it is faster in same-partition queries; 4) N-TS-HP has faster index construction, smaller index size, faster query and update than G-Tree by leveraging TD for overlay index construction; 5) Sketch can scale up to very large graphs (the only algorithm that can be applied on graph FR), but its query efficiency is much slower than N-PC-CP and P-PT-CP due to the light index.
@@ 7.3. PSP Application Guidance
Based on the experimental analysis above, we clarify the misconceptions mentioned in Section 1 and conclude the following observations: 1) All PSP indexes are essentially interrelated because they all consist of three dimensions in our PSP index scheme: graph partition method, PSP strategy, and shortest path algorithm. Moreover, we could combine the elements of these three dimensions to create a new PSP index; 2) It is important to choose the right partition structure for different types of graphs. Generally, for road networks, planar and hierarchical PSP indexes are more suitable than core-periphery PSP due to their efficient query and index update efficiency. On the other hand, core-periphery PSP indexes are more suitable for complex networks due to their excellent scalability; 3) Although PSP indexes typically have faster indexing efficiency, smaller index size, and better scalability than their non-partitioned counterparts, they sacrifice query and index update efficiency.
Figure 18. Our Decision Tree for PSP Index Design
@@
8. Future Directions
The insightful structure-based partition classification and universal PSP scheme proposed in this paper provide a systematical way to analyze and design the PSP index. Besides, the novel no-boundary and post-boundary strategies significantly enrich the design space for the PSP index. However, there are still some unresolved problems with the PSP index. Next, we will discuss some future directions and open challenges to inspire researchers in this field.
1) I/O-efficient and Distributed PSP Algorithms. There are two general ways to scale up graph algorithms: scale up to external memory on a single machine (out-of-core PSP index), or scale out to a distributed cluster (distributed PSP index). Since the PSP indexes naturally decompose a graph into several subgraphs, we could regard each subgraph (or its partition index) as a whole for disk accessory to reduce the I/O numbers or distribute each subgraph to a server in the cluster to reduce communication. Nevertheless, how to partition the graph for fewer I/O numbers or communication workload according to the network structure and algorithm procedure is still an open challenge. Moreover, how to optimize the PSP strategy and underlying SP algorithm simultaneously according to the I/O accessory or communication pattern is also an open challenge.
2) Automatic PSP Index Design System. Although we have propose a universal PSP index scheme and insightful guideline for the PSP index design, they are essentially experience-based. In the era of Artificial Intelligence (AI), how to leverage AI to automatically design the PSP index according to the system requirements is an open challenge. Moreover, how to make the elements of different dimensions in PSP index scheme portable such that the system could react promptly with the evolving system requirements is also an open challenge.
3) High Throughput System Based on PSP Index. Another future direction is to leverage PSP index for high throughput shortest path query processing. For example, road networks generally incur high throughput queries while the edge weight of road networks (travel time) could dynamically changes due to evolving traffic conditions. Therefore, how to achieve fast index maintenance and query processing simultaneously is important to enhance the system throughput. The PSP index could leverage parallelization among partition indexes to speed up the index update. Besides, adopting 2-hop labeling as the underlying SP algorithm also contributes to a fast query processing, making the PSP index a promising option for building path-finding system with high throughput. Nevertheless, how to design such a PSP index for high throughput query processing on dynamic road networks is an open challenge. Besides, partial edge updates in a partition may not influence all queries on other partitions. Thus how to identify the affected queries such that the system could still leverage the unaffected index for fast query processing is another open challenge.
@@ 9. Conclusions
In this work, we systematically study the partitioned shortest path index and propose a universal PSP scheme by considering three dimensions: SP algorithm, partition method, and PSP strategy. We first briefly review the SP algorithms used in the PSP index and then propose a new structure-based classification to facilitate the selection of the partition method and conduct a comprehensive review of corresponding partition methods and PSP indexes. For the PSP strategy, we summarize the classic pre-boundary strategy and then propose two new strategies and pruning-based overlay graph optimization to improve the index performance. We also provide index maintenance methods for all PSP strategies. Furthermore, we construct our PSP index scheme by seamlessly coupling those three dimensions. We also propose five new PSP indexes that are either more efficient in query or update than the current state-of-the-art. Systematic and extensive experimental evaluations demonstrate the effectiveness of our proposed techniques and new indexes guided by the insightful PSP index scheme. An insightful PSP index design guideline is provided to help future researchers and practitioners design their indexes. Lastly, the future directions and open challenges for the PSP index are also discussed.